[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun Oct 08, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Oct 9 20:35:47 UTC 2017


HSU Shift Log for Sun Oct 08 2017 22:27:28 GMT-0400 (EDT)

Shift: R

Forecaster: R Berg

Technical Issues Summary:

ATCF issues:

N-AWIPS issues: ASCAT stopped showing up in N-AWIPS after the 16z ASCAT-B  
pass. The data are available on the ASCAT website.

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171009/b9f033c8/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list